How do you solve this equation? I think you have to multiply one of them by a negative and add them together, but I'm unsure. Please advise.
2x + 3y = 4
3x + 2y = 6

Found 2 solutions by ankor@dixie-net.com, checkley71:
Answer by ankor@dixie-net.com(22740)

2x + 3y = 4
3x + 2y = 6
:
Adding these will not help anything. We want the value of x or y in both equations
to be equal, they can have opposite signs, then we can add, or same signs we can subtract
:
Lets try to get y to have the same value
Mult equation 1 by 2, and equation 2 by 3:
:
4x + 6y = 8
9x + 6y = 18
-----------------subtracting will eliminate y, we can solve for x
-5x + 0y = -10
+5x = +10; multiplied both side by -1 to get rid of the negatives
x = 10/2
x = 5
:
Use the 1st original equation to find y. Substitute 5 for x
2(5) + 3y = 4
10 + 3y = 4
3y = 4 - 10
3y = -6
y = -6/3
y = -2
:
Use the 2nd original equation to confirm out solutions
3x + 2y = 6
3(5) + 2(-2) =
10 - 4 = 10 confirms our solution
:
:
Did this help you?

Answer by checkley71(8403)

MULTIPLY THE FIRST EQUATION BY 3 THEN MULTIPLY THE SECOND EQUATION BY -2 & ADD THEM.
6X+9Y=12
-6X-4Y=-12
------------------------------
5Y=0
Y=0 ANSWER.
6X+9*0=12
6X=12
X=12/6
X=2 ANSWER.
PROOF
3*2+2*0=6
6=6